A 29-year German government bond (bund) has a face value of €150 and a coupon rate of 4% paid annually. Assume that the interest rate (in euros) is equal to 7.90% per year. What is the bond's PV?
Answer #1

4Hv51LCws8MEwAAAAASUVORK5CYII=

PV factor workings

=(1-(1+r)^-n)/r

.=+(1-(1+7.9%)^(-29))/7.9%=11.2627

=1/(1+r)^n

=+1/(1+7.9%)^29=.1102